OpenStudy (anonymous):

Solve algebraically for x: 4-√2x-5) = 1

OpenStudy (anonymous):

4-sqrt(2x-5) = 1 -sqrt(2x-5) = -3 sqrt(2x-5) = 3 square both sides 2x-5 = 9 2x = 14 x=7 To test that this is correct, plug 7 for x into the original equation: 4-sqrt(2(7) -5)...does this = 1? 4-sqrt (14-5) 4-sqrt(9) 4-3 = 1 Correct!

OpenStudy (anonymous):

Ahhh I see where I went wrong haha, thanks!

OpenStudy (anonymous):

Not a problem!